the lenght of a rectangle is five more than the width.If the perimeter of the rectangle is 18 meters,find the dimensions of the rectangle?

Found 2 solutions by checkley77, l-mayfield:
Answer by checkley77(12844)

L=W+5
2L+2W=18
2(W+5)+2W=18
2W+10+2W=18
4W=18-10
4W=8
W=8/4
W=2 ANS. FOR THE WIDTH.
W=2+5=7 ANS. FOR THE LENGTH.
PROOF:
2*7+2*2=18
14+4=18
18=18

Answer by l-mayfield(2)

P=18 To find P you have 2L + 2W you know that L is equal to 5+W. Substitute the 5+W in the equation. 18=2L+2W => 18=2(5+W)+2W => 18=10+2W+2W => subtract 10 from both sides 18-10=10-10+2W+2W => 8=4W => divide both sides by 4 to get 2=W. You know that L=5+W so 5+2 = 7. L=7 and W=2. Test 2(7)+2(2)=14+4=18